# Schur’s Positive-Definite Network: Deep Learning in the SPD cone with structure

### v3

@@ 3.1 Algorithmic foundations
The expressivity of SpodNet comes from the flexibility to use any arbitrary functions $f$ and $g$ in the updates whilst guaranteeing that $\bm{\Theta}$ remains in the SPD cone. Namely, additional structure such as sparsity can trivially be imposed on $\bm{\Theta}$ through $f$. We next detail the computational cost of fully updating $\bm{\Theta}$ through one SpodNet layer.
@@ 3.2 Improving the update complexity
Identifying all blocks yields $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$ which can be computed in $\mathcal{O}(p^{2})$ if one has access to $\bm{W}$. This can be achieved using Equation 3, and involves only operations in $\mathcal{O}(p^{2})$. The property $\bm{\Theta}^{+}=[\bm{W}^{+}]^{-1}$ is satisfied by using the same inversion formula. ∎
@@ 4.1 Inferring sparse precision matrices
where $\|\bm{\Theta}\|_{1,\mathrm{off}}$ denote the off-diagonal $\ell_{1}$ norm of $\bm{\Theta}$, equal to $\sum_{i\neq j}|\Theta_{ij}|$. It can be efficiently computed (Rolfs et al., 2012; Mazumder and Hastie, 2012; Hsieh et al., 2014) but suffers from the known limitations from the model-based approaches (Adler and Öktem, 2018; Arridge et al., 2019). To circumvent these limitations, two learning-based approaches have been proposed. Belilovsky et al. (2017) introduced DeepGraph, a CNN that directly maps empirical covariance matrices to the graph structures (i.e. to the support of the precision matrix). We emphasize that this model does not estimate $\bm{\Theta}$ but only its support. Shrivastava et al. (2020) proposed GLAD, a neural architecture based on unrolling an alternating minimization algorithm for solving the GLasso. Based on a Lagrangian relaxation, GLAD iterates over two matrices $\bm{Z}$ and $\bm{\Theta}$ while learning step-size and thresholding parameters (see Section B.2 for more details). Out of these two matrices $\bm{Z}$ is sparse and $\bm{\Theta}$ is SPD, but GLAD fails to learn a single matrix that is both sparse and SPD. This limitation is highlighted in Figure 2 showing that, during training, $\bm{\Theta}$ and $\bm{Z}$ differ, $\bm{\Theta}$ is not sparse and $\bm{Z}$ is not SPD.
@@ 4.2 SpodNet for sparse SPD learning
Finally, for all models, we use for $g$ a small neural network that takes as input $\theta_{22},s_{22}$ and $(\bm{\theta}_{12}^{+})^{\top}[\bm{\Theta}_{11}]^{-1}\bm{\theta}_{12}^{+}$. Its positivity is ensured by using an absolute value function as final nonlinearity.
@@ 5 Experiments
We now illustrate SpodNet’s ability to learn SPD matrices with additional structure by using our three derived graph learning models (UBG, PNP and E2E) to learn sparse precision matrices, on both synthetic and real data. Additional details regarding the experimental setups can be found in Appendix A. PyTorch implementations are provided with the submission and will be made public upon acceptance.
@@ 5.2 Application to unsupervised graph learning on a real-world dataset
Figure 6: Graphs estimation on the Animals dataset with the GLasso, EGFM and our UBG model. The former and the latter yield cleaner and more interpretable clusters than the GLasso, but EGFM is specifically tailored for this task as opposed to our UBG model.
@@
We compare our approach with the GLasso (with the regularization parameter selected by cross-validation), and the Elliptical Graphical Factor Model (EGFM) proposed by Hippert-Ferrer et al. (2023). The latter is a Riemannian optimization method specifically designed to identify clusters within the data by estimating a precision matrix whose inverse is modeled as a low-rank matrix plus a positive diagonal. For a fair comparison, we adopt the same hyperparameters as outlined in the original paper, setting the rank and regularization parameter to 10.
Each method yields a precision matrix, which is used to construct a graph representing the connections between the $p$ animals. The graph is constructed by considering the absolute values of the precision matrix coefficients, which represent the strength of the connections, and by removing the diagonal elements to eliminate self-loops. The results are illustrated in Figure 6, where the nodes of the graphs are partitioned using the Louvain algorithm (Blondel et al., 2008). Qualitatively, we observe that UBG produces a coherent graph, with outcomes comparable to those of GLasso and EGFM. Additionally, the graph learned by our UBG model exhibits a cleaner structure compared to GLasso, with a sparser representation and finer clusters, such as the grouping of (Chimp, Gorilla). To quantitatively assess the quality of the obtained graphs, we calculate the modularity $m$ of the partitions: higher modularity values indicate better separation of the graph into distinct subcomponents (Newman, 2006). We find that the graph produced by UBG achieves a higher modularity ($m=0.78$) than GLasso ($m=0.61$) and slightly lower than EGFM ($m=0.86$). It is important to note that the EGFM is specifically designed to obtain clustered graphs, whereas our UBG model has no such inductive bias. These combined quantitative and qualitative results demonstrate that SpodNet generalizes effectively to unseen data and produces a coherent graph structure.
@@ 6 Conclusion
We have proposed Schur’s Positive-Definite Network (SpodNet), a novel learning module compatible with other standard architectures, offering strict guarantees of SPD outputs. The principal novelty of SpodNet comes from its ability to handle additional desirable structural constraints, such as elementwise sparsity which we used as an illutrative example throughout this paper. To the best of our knowledge, SpodNet layers are the first to offer strict guarantees of such highly non-trivial structure in the outputs. In future works, SpodNet could be leveraged to learn other additional structures beyond sparsity. We have shown how to leverage SpodNet to build neural architectures that outperform traditional methods in the context of sparse precision matrix estimation in various settings. Future directions of research include theoretical investigations in order to better understand the expressivity and derive convergence guarantees for our approach.
@@ 7 Reproducibility statement
We believe this paper fully discloses the information needed to reproduce the main experimental results. Our core contribution is the SpodNet layer: its general algorithmic framework is presented in Algorithm 1. Three specific models built using this framework for learning sparse and SPD matrices are introduced in Section 4.2, and the exact architectures used for each of them throughout our experiments are explained in detail in Section A.2. The conducted experiments are described in Section 5 with additional details in Appendix A. PyTorch implementations are provided along with the submission.
8 Acknowledgments

### v4

@@ 3.1 Algorithmic foundations
The expressivity of SpodNet comes from the flexibility to use any arbitrary functions $f$ and $g$ in the updates whilst guaranteeing that $\bm{\Theta}$ remains in the SPD cone. Namely, additional structure such as sparsity can trivially be imposed on $\bm{\Theta}$ through $f$. Broadly speaking, constraints related to the values of the off-diagonal entries of $\bm{\Theta}$ and that can be imposed on the outputs of a neural network can be controlled directly. We next detail the computational cost of fully updating $\bm{\Theta}$ through one SpodNet layer.
@@ 3.2 Improving the update complexity
Identifying all blocks (namely $g(\Theta)$ with $\theta_{22}-\theta_{21}[\Theta_{11}]^{-1}\theta_{12}$, $\bm{w}_{12}$ with $-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}$ and $w_{22}$ with $\frac{1}{g(\bm{\Theta})}$) yields $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$ which can be computed in $\mathcal{O}(p^{2})$ if one has access to $\bm{W}$. This can be achieved using Equation 3, and involves only operations in $\mathcal{O}(p^{2})$. The property $\bm{\Theta}^{+}=[\bm{W}^{+}]^{-1}$ is satisfied by using the same inversion formula. ∎
@@ 4.1 Inferring sparse precision matrices
where $\|\bm{\Theta}\|_{1,\mathrm{off}}$ denote the off-diagonal $\ell_{1}$ norm of $\bm{\Theta}$, equal to $\sum_{i\neq j}|\Theta_{ij}|$. The data fidelity term $-\log\det(\bm{\Theta})+\langle\bm{S},\bm{\Theta}\rangle$ is the negative log-likelihood under Gaussian assumption, while the $\ell_{1}$ penalty enforces sparsity. It can be efficiently computed (Rolfs et al., 2012; Mazumder and Hastie, 2012; Hsieh et al., 2014) but suffers from the known limitations from the model-based approaches (Adler and Öktem, 2018; Arridge et al., 2019). To circumvent these limitations, two learning-based approaches have been proposed. Belilovsky et al. (2017) introduced DeepGraph, a CNN that directly maps empirical covariance matrices to the graph structures (i.e. to the support of the precision matrix). We emphasize that this model does not estimate $\bm{\Theta}$ but only its support. Shrivastava et al. (2020) proposed GLAD, a neural architecture based on unrolling an alternating minimization algorithm for solving the GLasso. Based on a Lagrangian relaxation, GLAD iterates over two matrices $\bm{Z}$ and $\bm{\Theta}$ while learning step-size and thresholding parameters (see Section B.2 for more details). Out of these two matrices $\bm{Z}$ is sparse and $\bm{\Theta}$ is SPD, but GLAD fails to learn a single matrix that is both sparse and SPD. This limitation is highlighted in Figure 2 showing that, during training, $\bm{\Theta}$ and $\bm{Z}$ differ, $\bm{\Theta}$ is not sparse and $\bm{Z}$ is not SPD.
@@ 4.2 SpodNet for sparse SPD learning
Finally, for all models, we use for $g$ a small neural network that takes as input $\theta_{22},s_{22}$ and $(\bm{\theta}_{12}^{+})^{\top}[\bm{\Theta}_{11}]^{-1}\bm{\theta}_{12}^{+}$. Its positivity is ensured by using an absolute value function as final nonlinearity. These input features are based on the intuition that the network should exploit information from the current state of the diagonal of $\bm{\Theta}$, the diagonal of $\bm{S}$ which is closely related to the diagonal of the GLasso estimator at convergence, and Schur’s complement which is added to the output of $g$.
@@ 5 Experiments
We now illustrate SpodNet’s ability to learn SPD matrices with additional structure by using our three derived graph learning models (UBG, PNP and E2E) to learn sparse precision matrices, on both synthetic and real data. Additional details regarding the experimental setups can be found in Appendix A. PyTorch implementations can be found in our GitHub repository [^2] https://github.com/Perceptronium/SpodNet.
@@ 5.2 Application to unsupervised graph learning on a real-world dataset
Figure 6: Graph estimation on the Animals dataset with the GLasso, EGFM, GLAD and our UBG model. The last three yield cleaner and more interpretable clusters than the GLasso, but EGFM is specifically tailored for finding clusters as opposed to our UBG model or GLAD.
@@
We compare our approach with the GLasso (with the regularization parameter selected by cross-validation), GLAD (which is trained with the same data as our model), and the Elliptical Graphical Factor Model (EGFM) proposed by Hippert-Ferrer et al. (2023). The latter is a Riemannian optimization method specifically designed to identify clusters within the data by estimating a precision matrix whose inverse is modeled as a low-rank matrix plus a positive diagonal. For a fair comparison, we adopt the same hyperparameters as outlined in the original paper, setting the rank and regularization parameter to 10.
Each method yields a precision matrix, which is used to construct a graph representing the connections between the $p$ animals. The graph is constructed by considering the absolute values of the precision matrix coefficients, which represent the strength of the connections, and by removing the diagonal elements to eliminate self-loops. The results are illustrated in Figure 6, where the nodes of the graphs are partitioned using the Louvain algorithm (Blondel et al., 2008). Qualitatively, we observe that UBG produces a coherent graph, with outcomes comparable to those of the three baselines. Additionally, the graph learned by our UBG model exhibits a cleaner structure compared to GLasso, with a sparser representation and finer clusters, such as the grouping of (Chimp, Gorilla). To quantitatively assess the quality of the obtained graphs, we calculate the modularity $m$ of the partitions: higher modularity values indicate better separation of the graph into distinct subcomponents (Newman, 2006). We find that the graph produced by UBG achieves a higher modularity ($m=0.78$) than GLasso ($m=0.61$) and GLAD ($m=0.72$), and slightly lower than EGFM ($m=0.86$). It is important to note that the EGFM is specifically designed to obtain clustered graphs, whereas our UBG model has no such inductive bias. These combined quantitative and qualitative results demonstrate that SpodNet generalizes effectively to unseen data and produces a coherent graph structure.
@@ 6 Conclusion
We have proposed Schur’s Positive-Definite Network (SpodNet), a novel learning module compatible with other standard architectures, offering strict guarantees of SPD outputs. The principal novelty of SpodNet comes from its ability to handle additional desirable structural constraints, such as elementwise sparsity which we used as an illutrative example throughout this paper. To the best of our knowledge, SpodNet layers are the first to offer strict guarantees of such highly non-trivial structure in the outputs. In future works, SpodNet could be leveraged to learn other additional structures beyond sparsity. We have shown how to leverage SpodNet to build neural architectures that outperform traditional methods in the context of sparse precision matrix estimation in various settings. Future research will be dedicated to improving the computational cost of our framework, theoretical understanding of the eigenvalues’ dynamics during training and of SpodNet’s expressivity, and deriving formal convergence guarantees.
@@ 7 Reproducibility statement
We believe this paper fully discloses the information needed to reproduce the main experimental results. Our core contribution is the SpodNet layer: its general algorithmic framework is presented in Algorithm 1. Three specific models built using this framework for learning sparse and SPD matrices are introduced in Section 4.2, and the exact architectures used for each of them throughout our experiments are explained in detail in Section A.2. The conducted experiments are described in Section 5 with additional details in Appendix A. PyTorch implementations can be found in our GitHub repository.
8 Acknowledgements
@@ B.3 GLAD’s non-SPD outputs
Additionally, $\bm{Z}$ loses its sparsity if projected onto the SPD cone and becomes singular (since it is actually projected onto the closed cone of semi-definite matrices). Going the other way around and trying to sparsify $\bm{\Theta}$ instead, by thresholding its off-diagonal entries following the support found by $\bm{Z}$, breaks its positive-definiteness guarantee (Figure 9).
@@
Figure 9: First row: Projecting GLAD-$\bm{Z}$ onto the PSD cone. Second row: Thresholding GLAD-$\bm{\Theta}$ following the support of GLAD-$\bm{Z}$.